OpenStudy (anonymous):

The range of F (x) = 3 sin 2 x is -1≤F(x)≤1 -2≤F(x)≤2 -3≤F(x)≤3

OpenStudy (anonymous):

f(x) = Asin(kx) + c where c is the axis of the curve and A is the amplitude.

OpenStudy (anonymous):

so k would be the range? making it 2?

OpenStudy (anonymous):

c +- A would be the range.

OpenStudy (anonymous):

Which would make it -1 / 1? because 2-3=-1 & 3-2=1. Right?

OpenStudy (anonymous):

there's no c, meaning c is 0

OpenStudy (anonymous):

Oh okay, I see now. Thank you so much!!